I need help with a question, Find the value of a in the equation
5
--- =
a+3
3
---
a-2

I think the answer is a= 2 3/8
Am I right, Can someone let me know?

Answer by sabanasir(22)

5 3
---- = ----
a + 3 a -2
do the cross multiplication
5(a-2)= 3(a+3)
5a-10=3a+9
2a=19
a=19/2
a= 9 1/2